1.Two large isolated parallel plates, 1m apart, are given equal and opposite charges; the potential difference between them 12V. Points X and Y are situated at 0.25m and 0.5m from the positive plate.

a.What is the difference in potential between X and Y

b.What is the value of the ration: electric field at Y/ electric field at X

c.How much work is done in moving a charge of 2C from X to Y

26


1
Expert's answer
2021-06-24T12:00:14-0400

Gives

Potential difference (V)=12V

d=1m

"r_x=0.25m,r_y=0.5m"

X and Y point are same plate (positive plate)

Part (a)

Potential difference between point X and Y

"\u2206V=V_x-V_y=12-12=0V"

Part(b)

Electric field ratio

"\frac{E_y}{E_x}=\frac{\frac{KQ}{r_x^2}}{\frac{KQ}{r_y^2}}=\frac{r_y^2}{r_x^2}=\frac{0.5^2}{0.25^2}=\frac{4}{1}"

Part(c)

X and Ypoint are same plate


"W=Q∆V=Q(V_x-V_y)=2\times0 =0J"